# Multivariate spatial models for small area estimation of species-specific forest inventory parameters Jeffrey W. Doser1, Malcolm S. Itter2, Grant M. Domke3, Andrew O. Finley4, 5

### v1

@@ 3 Results
Estimates of species-specific aboveground biomass from the multivariate spatial model across the 1,306 counties in the southern US showed close correspondence to design-based direct estimates across all twenty species (Figure 2). The average correlation coefficient between the model-based estimates and design-based estimates across the 20 species was 0.85, ranging from 0.96 for slash pine to 0.69 for southern red oak. Correlations with the kNN-based estimates were similarly high (Table 1). County-level estimates for the most abundant (loblolly pine) and least abundant (American beech) species are shown in Figure 3, along with associated 95% credible intervals as a characterization of uncertainty in the estimates. Maps of estimates and their uncertainty for all 20 species are provided in Supplemental Information S1.
@@
The model-based county-level estimates were consistently more precise across the twenty species compared to the design-based estimates. Across all species, 91.5% of county-level biomass estimates had higher precision (in terms of the coefficient of variation) compared to the design-based estimates, with mockernut hickory (98.9%) having the largest percentage improvements and longleaf pine (68.9%) having the lowest percentage improvements (Figure 4). Note that precision comparisons are only possible in counties where a species was observed on at least one FIA plot, as the design-based estimates do not have standard errors for counties where a species was not observed (i.e., all observations are zero).
@@
Visualizing the relative efficiency of model-based county-level biomass estimates compared to the design-based direct estimates highlights the spatial extent of improvement the proposed model provides (Figures 5, 6). The multivariate spatial model led to widespread improvements across species in essentially the entirety of the southern US study region (Figure 6). Noticeably, the multivariate spatial model had inferior performance to the design-based estimates in the western portion of the region as well as in southern Florida. Similar patterns emanate in maps of relative efficiency for individual species (Supplemental Information S1). For example, live oak, which has a complex spatial distribution spanning much of the southern and eastern coast of the study region, has large precision gains across Florida and much of Georgia, while precision is slightly decreased in much of the western portion of its range, particularly in several isolated counties (Supplemental Information S1: Figure S26).
@@ 3.1 Simulation study
The proposed model-based estimator, the direct estimator, or the kNN-based estimator did not have any systematic bias, although variation in bias across counties tended to be larger for the model-based estimator than the direct estimator and kNN estimator (Supplemental Information S2: Figure S1). Importantly, the model-based estimator had consistently higher accuracy (lower RMSE) compared to both the direct estimator and the kNN estimator (Supplemental Information S2: Figure S2).
@@ 4 Discussion
While NFIs have long provided statistically reliable information on forest resources at national and other large spatial scales, there is increased demand for estimates within smaller spatial scales (Dettmann et al.,, 2022; Prisley et al.,, 2021). Standard design-based approaches are limited in their ability to provide reasonable precision of small area estimates, particularly when estimating forest parameters by species or species groups. Here we developed a multivariate spatial modeling framework to enable small area estimation of species-specific forest inventory parameters. In a simulation study, we found less bias and higher accuracy from our model compared to a design-based estimator and non-parametric kNN estimator. We applied our model to estimate county-level aboveground biomass for the 20 most abundant tree species in the southern US, which resulted in substantial precision gains relative to the design-based direct estimates. More broadly, our framework can be applied to estimate species-level forest parameters (e.g., biomass, carbon, volume) within management-relevant spatial scales to support diverse forest management objectives.
County-level estimates of species-specific aboveground biomass in the southern US showed high correspondence with design-based estimates (Figure 2) and substantial improvements in estimate precision across the 20 species and 1,306 counties (Figures 4, 5). Unlike the design-based direct estimator, the model-based approach can readily provide estimates for small areas without any inventory plots by leveraging information from auxiliary covariates and spatially proximate locations. Further, it is common for individual species to never be observed within the set of inventory plots in a given small area, particularly when the number of plot locations is small. While design-based estimates in such situations are always estimated at 0 with no corresponding standard error, the proposed model does not have such restrictions. The model-based estimator can readily provide uncertainty predictions in small areas where a species is never observed, and is also not forced to estimate the parameter at 0. This is particularly relevant for predicting attributes of rare species that might be missed in relatively small fixed area plots such as those used in FIA but occur in other portions of the small area of interest. In such a situation, the multivariate spatial model can use information from the auxiliary variables in the model, the spatial random effects, and the presence of other species to help estimate the forest parameter of interest.
@@
Estimates from our hierarchical Bayesian model were highly correlated with kNN-based estimates (Table 1). kNN methods are frequently applied for forest parameter mapping (McRoberts et al.,, 2007; McRoberts,, 2012), and their ability to handle multivariate and non-Gaussian data makes them an attractive solution for estimating species-specific forest parameters (Breidenbach et al.,, 2010). However, kNN estimates do not have valid second order statistical properties, which has resulted in uncertainty quantification of kNN-based maps being a notorious limitation. This is particularly problematic if small area estimates are used in downstream models that take biomass, carbon, or tree volume as inputs (e.g., economic or ecological process models). The hierarchical Bayesian approach described here generates a full posterior predictive distribution for each small area estimate, which readily enables full uncertainty propagation in resulting estimates that can be accounted for in subsequent analyses.
@@
Despite the advances our proposed model provides, there are certain limitations that require future consideration. First, the log transformation of biomass occasionally resulted in extremely large uncertainty for the small area estimates (e.g., western Texas for live oak and water oak, Supplemental Information S1: Figures S6 and S19). Given this only occurred on the boundaries of our study region, this could be related to limitations of the isotropic and stationary exponential covariance function used to estimate the spatial processes. The use of more complex covariance functions (e.g., anisotropic functions) could help mitigate this pattern, which have been shown to provide improvements over isotropic functions in ecological applications (e.g., Finley, 2011). Additionally, counties along the edge of a species distribution within the study area often had more minimal precision gains relative to counties within the core of its range (e.g., laurel oak, live oak, slash pine). This could similarly reflect limitations of the isotropic exponential covariance function, or could result from localized drivers of biomass at these range edges that could not be explained by the covariates included in the analysis. Finally, care must be used when determining the size of the grid needed for generating the small area estimates (Finley et al.,, 2024). If the grid is not fine enough to represent spatial variation adequately throughout the small area of interest, the resulting estimates can be unstable. To ensure small area estimates are stable, we recommend predicting across multiple grid resolutions for at least a subset of the small areas of interest and comparing the resulting estimates. Because increasing the resolution of the grid will also increase the computational burden of the prediction task, one should choose the lowest resolution grid at which the estimates stabilize upon a single value.
We envision numerous extensions to our case study as well as the overall modeling framework. While we utilized wall-to-wall tree canopy cover data in combination with high-resolution climate data to aid in the prediction of species-level biomass across the southern US, incorporation of additional covariates related to forest productivity (e.g., soil characteristics) as well as Global Ecosystem Dynamics Investigation (GEDI) LiDAR data (Dubayah et al.,, 2020) could further improve species-level estimates. Future work could seek to develop a web-based interactive tool that allows users to directly access the model output and subsequently generate small area estimates within user-defined areas of interest. Finally, extending the model to a spatio-temporal context via a dynamic linear modeling framework (Gelfand et al., 2005a, ) could allow for improved insights on species-specific forest biomass and carbon change.
@@
NFIs are increasingly being called upon to provide insights on forest resources at small spatial scales. Forest parameter estimates by species are often desired at these scales to inform wildlife habitat protection, assessments of timber supply and products, and strategic management planning. Given the sample sizes that are feasibly collected by NFIs, classic design-based estimates are not able to provides estimates with reliable precision to meet these needs. The proposed multivariate spatial model developed here provides more precise estimates of species-specific forest parameters from NFIs, ultimately generating more opportunities for managers, landowners, and researchers to address small-scale management and research questions with NFI data.
@@ 5 Author Contributions
Jeffrey W. Doser: Writing – review & editing, Writing – original draft, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. Malcolm S. Itter: Writing – review & editing, Data curation, Methodology. Grant M. Domke: Writing – review & editing, Validation. Andrew O. Finley: Writing – review & editing, Methodology, Validation, Supervision, Resources, Funding acquisition.
@@ 6 Acknowledgments
We acknowledge the computing resources provided by North Carolina State University High Performance Computing Services Core Facility (RRID:SCR_022168). This work was supported by USDA Forest Service (FS), USDA FS/NCASI Partnership on Small Area Estimation, National Science Foundation (NSF) DEB-2213565 and DEB-1946007, and National Aeronautics and Space Administration (NASA) CMS Hayes-2023. The findings and conclusions in this publication are those of the authors and should not be construed to represent any official USDA or US Government determination or policy.
@@ 8 Data Availability Statement
All code and data associated with this manuscript are available on GitHub (https://github.com/doserjef/Doser_et_al_2025_SAE and will be posted on Zenodo upon acceptance.

### v2

@@ 3 Results
Estimates of species-specific aboveground biomass from the multivariate spatial model across the 1,306 counties in the southern US showed close correspondence to design-based direct estimates across all twenty species (Figure 2). The average correlation coefficient between the model-based estimates and design-based estimates across the 20 species was 0.85, ranging from 0.96 for slash pine to 0.69 for southern red oak (Table 1). County-level estimates for the most abundant (loblolly pine) and least abundant (American beech) species are shown in Figure 3, along with associated 95% credible intervals to characterize uncertainty in the estimates. Maps of estimates and their uncertainty for all 20 species are provided in Supplemental Information S1.
Despite the correspondence between model-based and design-based estimates, model-based estimates had a slight positive bias in county-level biomass estimates relative to the design-based estimates (Table 1). Patterns in bias were often species-specific, with multiple species showing increases in bias as sample size increased (e.g., water oak, laurel oak, slash pine; Figure 4). Residual assessments revealed moderate deviations from normality in the form of a heavy left tail (Supplemental Information S3) for many species, with the deviations being the most substantial for species with the largest amount of positive bias in county-level estimates. In general, the model-based estimates tended to provide more spatially smooth estimates relative to fine-scale heterogeneity represented in the direct estimates, as would be expected given the sharing of information between plots across space via the spatial autocorrelation structure and auxiliary variables in the model-based estimates (Supplemental Information S4). Large differences between model-based and design-based estimates occurred as a result of: (1) outlier counties within the species primary range that had a very high direct estimate in which the model provided substantially lower estimates more in line with surrounding counties (e.g., Figure 5a,b); (2) locations with high biomass where the model-based estimates tended to extend estimates of high biomass farther in space than the direct estimates (e.g., Figure 5c,d); (3) spatially disparate counties with positive biomass according to the direct estimate but often substantially lower biomass predicted from the model-based estimate. These patterns of correspondence between model-based and design-based estimates with slight positive bias in the model-based estimates were consistent when considering the entire data set as well as the hold-out assessments from the four-fold cross-validation (Table 2).
@@
The model-based county-level estimates were consistently more precise across the twenty species compared to the design-based estimates. Across all species, 91.5% of county-level biomass estimates had higher precision (in terms of the coefficient of variation) compared to the design-based estimates, with mockernut hickory (98.9%) having the largest percentage improvements and longleaf pine (68.9%) having the lowest percentage improvements (Table 1). Note that precision comparisons are only possible in counties where a species was observed on at least one FIA plot, as the design-based estimates do not have standard errors for counties where a species was not observed (i.e., all observations are zero).
@@
Visualizing the relative efficiency of model-based county-level biomass estimates compared to the design-based direct estimates highlights the spatial extent of improvement the proposed model provides (Figures 6, 7). The multivariate spatial model led to widespread improvements across species in essentially the entirety of the southern US study region (Figure 7). Noticeably, the multivariate spatial model had inferior performance to the design-based estimates in the western portion of the region as well as in southern Florida. Similar patterns emanate in maps of relative efficiency for individual species (Supplemental Information S1). For example, live oak, which has a complex spatial distribution spanning much of the southern and eastern coast of the study region, has large precision gains across Florida and much of Georgia, while precision is slightly decreased in much of the western portion of its range, particularly in several isolated counties (Supplemental Information S1: Figure S26).
@@ 3.1 Simulation study
The proposed model-based estimator and the direct estimator did not have any clear directional bias across the ten species, with the sign of the bias varying across the ten species (Supplemental Information S2: Table 1). Variation in bias across counties tended to be larger for the model-based estimator than the direct estimator (Supplemental Information S2: Figure S1). Importantly, the model-based estimator had consistently higher accuracy (lower RMSE) compared to the direct estimator (Supplemental Information S2: Figure S2).
@@ 4 Discussion
While NFIs have long provided statistically reliable information on forest resources at national and other large spatial scales, there is increased demand for estimates within smaller spatial scales (Dettmann et al.,, 2022; Prisley et al.,, 2021). Standard design-based approaches are limited in their ability to provide reasonable precision of small area estimates, particularly when estimating forest parameters by species or species groups. Here we developed a multivariate spatial modeling framework to enable small area estimation of species-specific forest inventory parameters. In a simulation study, we found less bias and higher accuracy from our model compared to a design-based estimator. We applied our model to estimate county-level aboveground biomass for the 20 most abundant tree species in the southern US, which resulted in substantial precision gains relative to the design-based direct estimates. More broadly, our framework can be applied to estimate species-level forest parameters (e.g., biomass, carbon, volume) within management-relevant spatial scales to support diverse forest management objectives.
County-level estimates of species-specific aboveground biomass in the southern US showed high correspondence with design-based estimates (Figure 2) and substantial improvements in estimate precision across the 20 species and 1,306 counties (Table 1, Figures 6, 7). Unlike the design-based direct estimator, the model-based approach can readily provide estimates for small areas without any inventory plots by leveraging information from auxiliary covariates and spatially proximate locations. Further, it is common for individual species to never be observed within the set of inventory plots in a given small area, particularly when the number of plot locations is small. While design-based estimates in such situations are always estimated at 0 with no corresponding standard error, the proposed model does not have such restrictions. The model-based estimator can readily provide uncertainty predictions in small areas where a species is never observed, and is also not forced to estimate the parameter at 0. This is particularly relevant for predicting attributes of rare species that might be missed in relatively small fixed area plots such as those used in FIA but occur in other portions of the small area of interest. In such a situation, the multivariate spatial model can use information from the auxiliary variables in the model, the spatial random effects, and the presence of other species to help estimate the forest parameter of interest.
@@
Despite the large improvements in estimate precision, the model-based estimates did consistently predict higher biomass relative to the design-based estimates, with the degree of positive bias varying substantially across species (Table 1, Figure 4). Bias was largest for loblolly pine, the most common and abundant species in the region (Figure 5c,d). Overestimation was fairly consistent within the core of loblolly pine’s range, with the largest areas of overestimation occurring in regions of high loblolly pine biomass. One potential reason for this is the limited ability of the covariates in Stage 2 of the model to predict areas of high biomass for loblolly pine, which results in the spatial component of the model attempting to explain such variation and providing more smooth estimates of high biomass in the surrounding region. This phenomenon is likely the main driver of the general tendency across all species to have, on average, higher model-based estimates compared to the direct estimates. Biomass of loblolly pine and other managed species in the study area (e.g., slash pine, longleaf pine) is heavily driven by management practices that are localized in space, and thus the ability to predict fine-scale biomass variation within a model-based framework may be limited for intensively managed species, unless such information can be incorporated into the model via additional auxiliary variables. More generally across all species, inclusion of covariates that characterize forest productivity (e.g., soil characteristics, disturbance history) as well as publicly available LiDAR data may provide more accurate estimates, particularly if covariates can explain fine-scale variation not accounted for by broad climatic and environmental variables. Additionally, use of the true FIA plot coordinates instead of the fuzzed plot coordinates used in this analysis would likely improve associations between species-specific biomass and the auxiliary variables, resulting in more accurate small area estimates. Such tasks represent important next steps to improve species-specific biomass estimates from FIA data such as those presented here.
While the higher estimates from the proposed model relative to the design-based estimates indicate areas of improvement for the model in the FIA case study, direct estimates of species-specific county-level biomass themselves may not accurately represent true characteristics of the county. When sample sizes within counties are small, the direct estimate solely relies on the small set of plots to generate an estimate (i.e., the sample mean). Sampling theory dictates that any individual direct estimate based on a small sample size may not itself be particularly close to the true value (i.e., given the large standard error of such an estimate). In the southern US case study, 25% of the counties contained less than 17 FIA plots, and thus such estimates could be substantially far from the true biomass density in the county. Thus, large discrepancies between model-based and design-based estimates in counties with low sample sizes does not inherently suggest inferior performance of the model-based approach in these areas, but rather represents a discrepancy between two estimates of some unknown, true population value. This concept complicates comparisons of accuracy between design-based and model-based estimates. Simulation studies such as the one implemented here provide an alternative approach for comparing model-based and direct estimators where the true underlying population is known.
@@
Despite the advances our proposed model provides in estimate precision, there are certain limitations that require future consideration. First, the log transformation of biomass occasionally resulted in extremely large uncertainty for the small area estimates (e.g., western Texas for live oak and water oak, Supplemental Information S1: Figures S6 and S19). Given this only occurred on the boundaries of our study region, this could be related to limitations of the isotropic and stationary exponential covariance function used to estimate the spatial processes. The use of more complex covariance functions (e.g., anisotropic functions) could help mitigate this pattern, which have been shown to provide improvements over isotropic functions in ecological applications (e.g., Finley, 2011). Additionally, counties along the edge of a species distribution within the study area often had more minimal precision gains relative to counties within the core of its range (e.g., laurel oak, live oak, slash pine). This could similarly reflect limitations of the isotropic exponential covariance function, or could result from localized drivers of biomass at these range edges that could not be explained by the covariates included in the analysis. Finally, care must be used when determining the size of the grid needed for generating the small area estimates (Finley et al.,, 2024). If the grid is not fine enough to represent spatial variation adequately throughout the small area of interest, the resulting estimates can be unstable. To ensure small area estimates are stable, we recommend predicting across multiple grid resolutions for at least a subset of the small areas of interest and comparing the resulting estimates. Because increasing the resolution of the grid will also increase the computational burden of the prediction task, one should choose the lowest resolution grid at which the estimates stabilize upon a single value.
@@
NFIs are increasingly being called upon to provide insights on forest resources at small spatial scales. Forest parameter estimates by species are often desired at these scales to inform wildlife habitat protection, assessments of timber supply and products, and strategic management planning. Given the sample sizes that are feasibly collected by NFIs, classic design-based estimates are not able to provide sufficient precision to meet these needs. The proposed multivariate spatial model developed here provides more precise estimates of species-specific forest parameters from NFIs, ultimately generating more opportunities for managers, landowners, and researchers to address small-scale management and research questions with NFI data.
@@ 5 Author Contributions
Jeffrey W. Doser: Writing – review & editing, Writing – original draft, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. Malcolm S. Itter: Writing – review & editing, Data curation, Methodology. Grant M. Domke: Writing – review & editing, Validation. Andrew O. Finley: Writing – review & editing, Methodology, Validation, Conceptualization, Supervision, Resources, Funding acquisition.
@@ 6 Acknowledgments
We thank two anonymous reviewers for their insightful comments that improved the manuscript. We acknowledge the computing resources provided by North Carolina State University High Performance Computing Services Core Facility (RRID:SCR_022168). This work was supported by USDA Forest Service (FS), USDA FS/NCASI Partnership on Small Area Estimation, National Science Foundation (NSF) DEB-2213565 and DEB-1946007, and National Aeronautics and Space Administration (NASA) CMS Hayes-2023. The findings and conclusions in this publication are those of the authors and should not be construed to represent any official USDA or US Government determination or policy.
@@ 8 Data Availability Statement
All code and data associated with this manuscript are available on GitHub (https://github.com/doserjef/Doser_et_al_2025_SAE) and will be posted on Zenodo upon acceptance.